# The Geometry of Numerical Reasoning: Language Models Compare Numeric Properties in Linear Subspaces

### v1

@@ Abstract
This paper investigates whether large language models (LLMs) utilize numerical attributes encoded in a low-dimensional subspace of the embedding space when answering logical comparison questions (e.g., Was Cristiano born before Messi?). We first identified, using partial least square regression, these subspaces, which effectively encode the numerical attributes associated with the entities in comparison prompts. Further, we demonstrate causality, by intervening in these subspaces to manipulate hidden states, thereby altering the LLM’s comparison outcomes. Experimental results demonstrated that our results stand for different numerical attributes, which indicates that LLMs utilize the linearly encoded information for numerical reasoning.
@@
Ahmed Oumar El-Shangiti1       Tatsuya Hiraoka1       Hilal AlQuabeh1 Benjamin Heinzerling2,3       Kentaro Inui1,2,3 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) 2Tohoku University 3RIKEN ahmed.oumar@mbzuai.ac.ae
@@ 1 Introduction
Language models (LMs) store large amounts of world knowledge in their parameters Petroni et al. (2019); Jiang et al. (2020); Roberts et al. (2020); Heinzerling and Inui (2021); Kassner et al. (2021). While prior work has evaluated parametric knowledge mainly via behavioral benchmarks, more recent work has analyzed how knowledge is represented in activation space, for example, localizing relational knowledge to specific layers and token representations Meng et al. (2022); Geva et al. (2023); Merullo et al. (2024) or identifying subspaces that encode numeric properties such as an entity’s birth year Heinzerling and Inui (2024). However, analysis of LM-internal knowledge representation has been limited to simple factual recall, e.g, for queries like “When was Cristiano born?” (Answer: 1985) or “When was Messi born?” (Answer: 1987). If and how the mechanisms responsible for simple factual recall also participate in more complex queries, e.g., “Is Cristiano older than Messi?”, is not understood so far. A possible mechanism by which an LM answers this query is a multi-step process consisting of first recalling the respective birth years of the two entities, comparing the two years, and then selecting a corresponding answer.
@@
Figure 1: Summary of our approach. We extract contextualized numeric attribute activations and then train $k$-components PLS model on the activations to predict their values and then use the first component of the PLS model to do intervention at the last token of the second entity in the logical comparison.
In this study, we tackle the research question: do LLMs leverage the linear subspace of entity-numerical attributes when solving numerical reasoning tasks? We investigate whether the linear subspace is indeed used in the logical reasoning tasks. We first show the LLMs’ capability to solve the numerical reasoning tasks from the viewpoint of behavioral observation: testing the performance of the reasoning task with in-context learning (§3). We then look into the representation of LLMs (§4). We identify the linear subspace corresponding to numerical attributes with partial least-squares (PLS) and intervene in the representation to test whether the model utilizes the linearly represented information (see Figure 1).
The experimental results on the three numerical properties (the birth/death year of person and the latitude of location) demonstrate that LLMs leverage the numerical information represented in the linear subspace for the reasoning tasks.
@@ 2 Outline of Experiments
What is the year of death for Isaac Newton? 1727
@@
Cairo’s latitude higher than Jerusalem’s? Yes
@@ 2.1 Model and Dataset
In this work, we focus on the three numerical properties: the birth years of person entities, the death years of person entities, and the latitudes of location entities. Table 1 exemplifies the questions and expected responses for both tasks. For the knowledge extraction task, we create the question-answer pairs by extracting 5,000 entities alongside their numerical attributes from Wikidata Vrandečić and Krötzsch (2014). After filtering out entities that the LLM does not know (§3.1), we created the 5,000 questions about numerical reasoning that include two entities each. For all experiments, we used Llama3-8B-instruction following model (Dubey et al., 2024) as the LLM.
@@ 2.2 Design of Experiments
We conducted the experiments into two phases to investigate the LLM’s ability to utilize the linear subspace for numerical reasoning.
@@
In the second phase, we examined the inner workings of the LLM when solving the knowledge extraction (§4.1) and the numerical reasoning (§4.2). Here, we focus on analyzing the hidden state of each entity representation at a particular layer for knowledge extraction. For the case of numerical reasoning, we investigated the activations of the last token’s representation. We denote the hidden state of the $i$-th input at the $l$-th layer as $h^{(l)}_{i}$. To investigate whether knowledge of numerical attributes is stored in low-dimensional subspaces, we applied PLS for each representation Heinzerling and Inui (2024). Through PLS, we identified components representing the linear structure of each numerical attribute, allowing us to analyze how the LLM might utilize these subspaces for reasoning. To further test this, we intervened in the hidden state $h^{(l)}_{i}$ by incorporating the 1st PLS component $\vec{v}$, as follows:
@@
where $\alpha$ is a hyperparameter derived from the first PLS component, and $||v||$ is the Euclidian norm (L2-norm) of the vector $v$. Intuitively, this intervention edits the numerical attribute captured by the LLM. For instance, if the numerical information (Cristiano, born-in, 1985) is shifted to (Cristiano, born-in, 2020),a LLM that genuinely relies on a linear subspace for reasoning would adjust its interpretation accordingly, reflecting the change in its responses (Figure 1).
@@ 4 Internal Representation Experiments
(a) Birth Attribute
(b) Death Attribute
(c) Latitude Attribute
Figure 4: The effect of intervention in the identified subspace in each layer, compared to a random direction from a normal distribution.
The aim of this experiment is to train a PLS model to identify low-dimensional linear subspaces within the activation space, which could potentially be efficient in predicting numerical attributes for various entities. Then we fitted another PLS model to evaluate Yes/No comparison reasoning related to these numerical attributes. Finally, we demonstrated the causal relationship within these subspaces by implementing targeted interventions.
@@ 4.1 Prediction of numerical attributes with PLS
The training procedure consists of the following: (1) we first filter out the entities that the model predicted their comparison incorrectly (Section 3.2). (2) We feed a context vector that contains the comparison prompt (e.g., Was Cristiano born prior to Messi? (3) We extract the hidden states of the last token of each entity from the LLM’s hidden states at a particular layer. (4) These hidden states are then used to train a PLS model with a $5$ component to predict the corresponding numerical attribute of each entity based on their corresponding model representation (activations). Figure 2 depict the results achieved by $5$ components PLS model, measured by the coefficient of determination $R^{2}$. The goodness of fit exceeds $0.8$ for all measured properties, indicating that the information encoded in these attributes can be extracted with using low-dimensional (linear) subspaces.
@@
LC $53$ $56$ $50$ $37.8$ $55$ $51.2$ $55$ $50$ $50$ $50.2$
Table 2: Experiments 1 and 2’s Results for 3 tasks, and 10 different prompts for each. The accuracy of exact matching is reported, except for Latitute task, where we relaxed the predicted and ground truth to be rounded to the integer part. BP: Birth Prediction, DP: Death Prediction, LP: Latitude Prediction, BC: Birth Comparison, DC: Death Comparison, LC: Latitude Comparison
4.2 Logical Comparison with PLS
@@ 4.3 Intervention using PLS Components Vector
While the previous experiments with the PLS model establish correlation, they do not demonstrate causality. For this purpose, we perform interventions at a particular token within a designated model layer, chosen based on the correlation strength identified in predicting numerical attributes from each task (Section 4.1). We fix the first entity and intervene at the last token of the second entity. This token’s hidden state is then updated by a scaled version of the first component direction from the PLS model to the original hidden state $h^{(l)}_{i}$ as illustrated in equation 1.
In Figure 4 we compare the effect of our intervention per layer against a random vector from the normal distribution. The results clearly demonstrate the superiority of our intervention method, particularly evident in Subfigures $a$ and $b$. In subfigure $c$, related to the Latitude numeric attribute, the gap between our method and the baseline narrows, suggesting that the direction may not be significant for this attribute. This could reflect the mode’s nearly random response in the behavior experiment (Section 3.2). Additionally, the intervention’s effect is notable only in the first $\approx 50\%$ of the model layers, after which it diminishes to zero, aligned with prior research on inference time theory. We also tested the generalization of our approach on unseen samples, as shown in appendix 5.
@@ 5 Conclusion
In this research, we empirically demonstrate that the model answers numerical reasoning questions, such as "Was Cristiano born before Messi?" using a two-step process. First, it extracts numerical attributes for each entity from a linear subspace. The second step involves utilizing these linear directions to answer the logical question. Specifically, subspaces are identified through PLS regression, where directions in low-dimensional subspaces of the activation space encode numerical property information. We illustrate this approach using three numerical attributes: Birth, Death, and Latitude. The reasoning step is validated using causal interventions along the direction of the first component of the PLS model, where these interventions successfully alter the model’s answers.
@@ 6 Limitations
  • Single Model Evaluation: This study focuses solely on the performance of a single large language model, LLaMA-3 (8B-instruction). While this model provides valuable insights, further experiments involving other models, particularly those with different architectures or training regimes, would be necessary to generalize our conclusions.

### v2

@@ Abstract
This paper investigates whether large language models (LLMs) utilize numerical attributes encoded in a low-dimensional subspace of the embedding space when answering questions involving numeric comparisons, e.g., Was Cristiano born before Messi?. We first identified, using partial least squares regression, these subspaces, which effectively encode the numerical attributes associated with the entities in comparison prompts. Further, we demonstrate causality, by intervening in these subspaces to manipulate hidden states, thereby altering the LLM’s comparison outcomes. Experiments conducted on three different LLMs showed that our results hold across different numerical attributes, indicating that LLMs utilize the linearly encoded information for numerical reasoning.
@@
Ahmed Oumar El-Shangiti1       Tatsuya Hiraoka1       Hilal AlQuabeh1 Benjamin Heinzerling3,2       Kentaro Inui1,2,3 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) 2Tohoku University 3RIKEN ahmed.oumar@mbzuai.ac.ae
@@ 1 Introduction
Language models (LMs) store large amounts of world knowledge in their parameters Petroni et al. (2019); Jiang et al. (2020); Roberts et al. (2020); Heinzerling and Inui (2021); Kassner et al. (2021). While prior work has evaluated parametric knowledge mainly via behavioral benchmarks, more recent work has analyzed how knowledge is represented in activation space, for example, localizing relational knowledge to specific layers and token representations Meng et al. (2022); Geva et al. (2023); Merullo et al. (2024) or identifying subspaces that encode numeric properties such as an entity’s birth year Heinzerling and Inui (2024). However, analysis of LM-internal knowledge representation has been limited to simple factual recall, e.g., for queries like “When was Cristiano born?” (Answer: 1985) or “When was Messi born?” (Answer: 1987). If and how the mechanisms responsible for simple factual recall also participate in more complex queries, e.g., “Is Cristiano older than Messi?”, is not understood so far. A possible mechanism by which an LLM answers this query is a multi-step process consisting of first recalling the respective birth years of the two entities, comparing the two years, and then selecting a corresponding answer.
@@
Figure 1: Summary of our approach. We extract contextualized numeric attribute activations and then train $k$-components PLS model on the activations to predict their values and then use the first component of the PLS model to do an intervention at the last token of the second entity in the logical comparison.
In this study, we tackle the research question: do LLMs leverage the linear subspace of entity-numerical attributes when solving numerical reasoning tasks? We investigate whether the linear subspace is indeed used in the logical reasoning tasks. We first show the LLMs’ capability to solve the numerical reasoning tasks from the viewpoint of behavioral observation: testing the performance of the reasoning task with in-context learning (§3). We then examine the representations of LLMs (§4). We identify the linear subspace corresponding to the numerical attributes with partial least-squares (PLS Wold et al. (2001)) and intervene in the representation to test whether the model utilizes the linearly represented information (see Figure 1).
The experimental results on the three numerical properties (the birth/death year of a person and the latitude of location) and on three LLMs (LLama3 8B Dubey et al. (2024), Mistral 7B Jiang et al. (2023), and Qwen2.5 7B Team (2024) all instruction based models) demonstrate that LLMs leverage the numerical information represented in the linear subspace for the reasoning tasks.
@@ 2 Outline of Experiments
What is Isaac Newton’s year of death? 1727
@@
Is Cairo’s latitude higher than Jerusalem’s? Yes
@@ 2.1 Model and Dataset
In this work, we focus on the three numerical properties: the birth years of person entities, the death years of person entities, and the latitudes of location entities. Table 1 exemplifies the questions and expected responses for both tasks. For the knowledge extraction task, we create the question-answer pairs by extracting 5,000 entities alongside their numerical attributes from Wikidata Vrandečić and Krötzsch (2014). After filtering out entities that the LLM does not know (§3.1), we created the 5,000 questions about numerical reasoning that include two entities each. For all experiments, we used Llama3-8B-instruction following model (Dubey et al., 2024) as the LLM and later validate our finding on two additional models (see § 4.3).
@@ 2.2 Design of Experiments
We conducted the experiments in two phases to investigate the LLM’s ability to utilize the linear subspace for numerical reasoning.
@@
In the second phase, we examined the inner workings of the LLM when solving the knowledge extraction (§4.1) and the numerical reasoning (§B.1). Here, we focus on analyzing the hidden state of each entity representation at a particular layer for knowledge extraction. For the case of numerical reasoning, we investigated the activations of the last token’s representation. We denote the hidden state of the $i$-th input at the $l$-th layer as $h^{(l)}_{i}$. To investigate whether knowledge of numerical attributes is stored in low-dimensional subspaces, we applied PLS Wold et al. (2001) for each representation Heinzerling and Inui (2024). Partial Least Squares (PLS) offers an alternative to Principal Component Analysis (PCA) for dimensionality reduction, especially when predicting one set of variables from another. PLS seeks to maximize the covariance between the input matrix X and the response matrix Y by projecting both onto a latent space. Through PLS, we identified components that represent the linear structure of each numerical attribute, allowing us to analyze how the LLM might utilize these subspaces for reasoning. To further test this, we intervened in the hidden state $h^{(l)}_{i}$ by incorporating the 1st PLS component $v$, as follows:
@@
where $\alpha$ is a hyperparameter derived from the first PLS component, and $||v||$ is the Euclidian norm (L2-norm) of the vector $v$. Intuitively, this intervention edits the numerical attribute captured by the LLM. For instance, if the numerical information (Cristiano, born-in, 1985) is shifted to (Cristiano, born-in, 2020), an LLM that genuinely relies on a linear subspace for reasoning would adjust its interpretation accordingly, reflecting the change in its responses (Figure 1).
@@ 4 Internal Representation Experiments
Figure 3: The effect of the intervention—specifically, the ratio of flipped answers after performing intervention—was analyzed within the identified model subspace of each layer and compared to the effects observed in a randomly selected direction sampled from a normal distribution.
This experiment aims to train a PLS model to identify low-dimensional linear subspaces within the activation space, which could potentially be efficient in predicting numerical attributes for various entities. We then demonstrated the causal relationship within these subspaces by implementing targeted interventions which shows that indeed there is a causal effect between the identified linear subspaces and the logical comparison answers by the model. We validate our hypothesis by running three models on three numerical attributes.
We also fitted another PLS model to evaluate Yes/No comparison reasoning related to these numerical attributes (see appendix  B.1).
@@ 4.1 Prediction of numerical attributes with PLS
The training procedure consists of the following steps: (1) we first filter out the entities that the model predicted their comparison incorrectly (Section 3.2). (2) We feed a context vector that contains the comparison prompt (e.g., Was Cristiano born prior to Messi?) (3) We extract the hidden states of the last token of each entity from the LLM’s hidden states at a particular layer. (4) These hidden states are then used to train a PLS model with a $5$ component to predict the corresponding numerical attribute of each entity based on their corresponding model representation (activations). Figure 2 depicts the results achieved by $five$ components PLS model, measured by the coefficient of determination $R^{2}$. The goodness of fit exceeds $0.8$ for all measured properties, indicating that the information encoded in these attributes can be extracted with low-dimensional (linear) subspaces.
@@
LC $53.0$ $56.0$ $50.0$ $37.8$ $55.0$ $51.2$ $55.0$ $50.0$ $50.0$ $50.2$
Table 2: Experiments 1 and 2’s Results for three tasks, and 10 different prompts for each. The accuracy of exact matching is reported, except for the Latitude task, where we relaxed the predicted and ground truth to be rounded to the integer part. BP: Birth Prediction, DP: Death Prediction, LP: Latitude Prediction, BC: Birth Comparison, DC: Death Comparison, LC: Latitude Comparison
@@ 4.2 Intervention using PLS Components Vector
While the previous experiments with the PLS model establish correlation, they do not demonstrate causality. For this purpose, we perform interventions at a particular token within a designated model layer, chosen based on the correlation strength identified in predicting numerical attributes from each task (Section 4.1). We fix the first entity and intervene at the last token of the second entity. This token’s hidden state is then updated by a scaled version of the first component direction from the PLS model to the original hidden state $h^{(l)}_{i}$ as illustrated in equation (1).
In Figure 3 we compare the effect of our intervention per layer against a random vector from the normal distribution. It is measured by the Effect of Intervention metric (EI) (equation 2), $f$ and $f^{\prime}$ are the clean and patched models.
$\text{EI}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\,[f(x_{i})\neq f^{\prime}(x_{i})]$ (2)
The results clearly demonstrate the superiority of our intervention method, particularly evident in Subfigures $a$ and $b$. In subfigure $c$, related to the Latitude numeric attribute, the gap between our method and the baseline narrows, suggesting that the direction may not be significant for this attribute. This could reflect the mode’s nearly random response in the behavior experiment (Section 3.2). Additionally, the intervention’s effect is notable only in the first $\approx 50\%$ of the model layers, after which it diminishes to zero, aligned with prior research on inference time theory. We also tested the generalization of our approach on unseen samples, as shown in appendix, Figure 9 and additional models (see § 4.3).
Figure 4: $R^{2}$ score of predicting entity’s birth years attributes, using a 5-Component PLS model trained on Mistral 7B Instruct activations.
Figure 5: $R^{2}$ score of predicting entity’s birth years attributes, using a 5-Component PLS model trained on Qwen2.5 7B Instruct activations.
Figure 6: The effect of the intervention(i.e. the ratio of the flipped answers) in the identified subspace in each layer of the Mistral 7B Instruct model, compared to a random direction from a normal distribution.
Figure 7: The effect of the intervention(i.e., the ratio of the flipped answer) in the identified subspace in each layer of the Qwen2.5 7B instruct model, compared to a random direction from a normal distribution.
Task Model Prompts
1 2 3 4 5 6 7 8 9 10
BP Mistral 7B 72.65 72.63 74.68 75.36 73.64 75.44 73.86 74.81 72.90 73.56
Qwen2.5 7B 40.82 34.68 33.95 34.59 33.95 36.72 36.96 32.61 39.07 34.32
BC Mistral 7B 53.60 64.84 64.02 53.10 61.88 57.66 53.00 67.06 64.68 50.00
Qwen2.5 7B 29.20 58.10 38.88 26.22 49.76 40.54 6.20 3.84 9.16 6.98
Table 3: Exact Matching Accuracy of Mistral 7B and Qwen2.5 7B Models on Birth Date Numerical property extraction and Comparison Tasks Across Prompt Variations. All models are instruction-based models. BP: Birth Prediction and BC: Birth Comparison tasks are evaluated.
4.3 Experiments on Additional Models
To further validate our hypothesis generalization, we run the same experiments on two additional language models for the birth property. Those additional models are Mistral-7B-intruct Jiang et al. (2023) and Qwen2.5-7B-Instruct Team (2024).
PLS models trained on models’ activation have crossed an $R^{2}$ score of $0.8$ suggesting that the information encoded in those models’ activations can be extracted using low-dimensional (linear) subspaces (see Figure 4 and Figure 5).
The Effect of Intervention (EI) results shown in Figures 6 and  7 of the Mistral 7B Instruct and Qwen-2.5 7B Instruct models, respectively, demonstrate the same behavior seen in the previous experiments. For the EI of Mistral, we can see that the peak was around the $11$th layer and then continued to decrease until it finally disappeared around the $16$ layer (Figure 6). When compared to other models, Qwen2.5 has shown two clear differences. First, we can observe two peaks for the EI with almost the same value of the EI, early around the $third$ layer and later one around layer $12$, while other models have shown only one peak. Second, unlike other models, Qwen2.5 7B kept bouncing around almost the same EI values and suddenly become None at around layer $16$ (Figure 7). One reason that might explain the difference between Qwen2.5 7B and other models, is that Qwen2.5 7B uses only $28$ layers, while other models in the experiments are formed of $32$ layers.
@@ 5 Conclusion
In this research, we empirically demonstrate that the model answers numerical reasoning questions, such as "Was Cristiano born before Messi?" using a two-step process. First, it extracts numerical attributes for each entity from a linear subspace. The second step involves utilizing these linear directions to answer the logical question. Specifically, subspaces are identified through PLS regression, where directions in low-dimensional subspaces of the activation space encode numerical property information. We illustrate this approach using three numerical attributes: Birth, Death, and Latitude across three LLMs. The reasoning step is validated using causal interventions along the direction of the first component of the PLS model, where these interventions successfully alter the model’s answers.
6 Ethical Statement
@@ Appendix B Related Work
B.1 Logical Comparison with PLS
(a) Birth Attribute
(b) Death Attribute
(c) Latitude Attribute